OpenStudy (anonymous):

Help pls If the quadratic formula is used to solve 2x2 = 8x - 3, what are the solutions?

OpenStudy (anonymous):

Just have to make 0 one side then use this: http://www.purplemath.com/modules/quadform.htm

OpenStudy (solomonzelman):

The solutions are same with any method.

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm 2x^2= 8x - 3}\) \(\large\color{blue}{ \rm 2x^2 - 8x =- 3}\) \(\large\color{blue}{ \rm 2x^2 - 8x +8=- 3+8}\) \(\large\color{blue}{ \rm 2x^2 - 8x +8=5}\) \(\large\color{blue}{ \rm 2(x^2 - 4x +4)=5}\) \(\large\color{blue}{ \rm (x^2 - 4x +4)=5/2}\) \(\large\color{blue}{ \rm (x-2)^2=5/2}\) \(\large\color{blue}{ \rm (x-2)^2=10/4}\) \(\large\color{blue}{ \rm x-2=±\sqrt{10}/2}\) \(\large\color{blue}{ \rm x=2±\sqrt{10}/2}\)